# Double Robust high dimensional alpha test for linear factor pricing model

### v1

@@ A.4 Proof of Theorem 3
$\mathbb{P}\left(\frac{T_{SS}}{\sqrt{2\mbox{tr}({\bf R}^{2})}}\leq x,T||\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}||^{2}_{\infty}\hat{\zeta}-2\log N+\log\log N\leq y\right)\rightarrow\Phi(x)\cdot\exp\left\{-\frac{1}{\sqrt{\pi}}e^{-y/2}\right\}.$ (10)
@@
$\mathbb{P}\left(\frac{T_{SS}}{\sqrt{2\mbox{tr}({\bf R}^{2})}}\leq x,T||\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}||^{2}_{\infty}\hat{\zeta}\leq u_{N}(y)\right)\rightarrow\Phi(x)\cdot\exp\left\{-\frac{1}{\sqrt{\pi}}e^{-y/2}\right\}.$
@@
$\mathbb{P}\left(\sqrt{\frac{1}{2\operatorname{tr}\left(\mathbf{R}^{2}\right)}}N\omega_{T}^{-1}\underset{t_{1}\neq t_{2}}{\sum\sum}h_{t_{1}}h_{t_{2}}{\bf U}_{t_{1}}^{\top}{\bf U}_{t_{2}}+o_{p}\left(1\right)\leq x,\right.$
$\left.\quad E(r_{t}^{2})\left\|T^{-1/2}\sum_{t=1}^{T}{\bf U}_{t}\right\|_{\infty}^{2}+O_{p}\left(L_{T,N}\right)\leq u_{N}(y)\right)$
@@
where $L_{T,N}=T^{-1/4}\log^{1/2}(TN)+N^{-(1/6\wedge\delta/2)}\log^{1/2}(TN)+T^{-1/2}(\log N)^{1/2}\log^{1/2}(TN)$. Then, we just need to prove (A.4).
From the Theorem 2 in Feng et al. (2022a), the equation (A.4) holds if ${\bf U}_{t}$ follows the normal distribution. We then investigate the non-normal case. Let $\boldsymbol{\xi}_{t}=$ ${\bf U}_{t}\in\mathcal{R}^{N},i=1,2,\cdots,T$. For $\boldsymbol{z}=\left(z_{1},\cdots,z_{q}\right)^{\top}\in\mathcal{R}^{q}$, we consider a smooth approximation of the maximum function, namely,
@@
$V\left(\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{T}\right)$ $=\beta^{-1}\log\left(\sum_{j=1}^{N}\exp\left(\beta\sqrt{\frac{E(r_{t}^{2})}{T}}\sum_{i=1}^{T}\boldsymbol{x}_{i,j}\right)\right).$
By setting $\beta=T^{1/8}\log(T)$, equation (A.4) is equivalent to
@@
Suppose $\left\{\boldsymbol{Y}_{1},\boldsymbol{Y}_{2},\cdots,\boldsymbol{Y}_{T}\right\}$ are sample from $N\left(0,\mathbb{E}{\bf U}_{1}^{\top}{\bf U}_{1}\right)$, and independent with ${\bf U}_{1},\cdots,{\bf U}_{T}$ or write as $\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{T}\right)$. The key idea is to show that: $\left(W\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{T}\right),V\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{T}\right)\right)$ has the same limiting distribution as $\left(W\left(\boldsymbol{Y}_{1},\cdots,\boldsymbol{Y}_{T}\right),V\left(\boldsymbol{Y}_{1},\cdots,\boldsymbol{Y}_{T}\right)\right)$.
@@
Let $l_{b}^{2}(\mathbb{R})$ denote the class of bounded functions with bounded and continuous derivatives up to order 3. It is known that a sequence of randon variables $\left\{Z_{n}\right\}_{n=1}^{\infty}$ converges weakly to a random variable $Z$ if and only if for every $f\in l_{b}^{3}(\mathbb{R}),\mathbb{E}\left(f\left(Z_{n}\right)\right)\rightarrow\mathbb{E}(f(Z))$. It suffices to show that:
$\mathbb{E}\left\{f\left(W\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{T}\right),V\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{T}\right)\right)\right\}-\mathbb{E}\left\{f\left(W\left(\boldsymbol{Y}_{1},\cdots,\boldsymbol{Y}_{T}\right),V\left(\boldsymbol{Y}_{1},\cdots,\boldsymbol{Y}_{T}\right)\right)\right\}\rightarrow 0,$
@@
for every $f\in l_{b}^{3}\left(\mathbb{R}^{2}\right)$ as $(N,T)\rightarrow\infty$. Because $h_{1},\dots,h_{T}$ with independent ${\bf U}_{1},\dots,{\bf U}_{T}$, we just need to prove
$\mathbb{E}\left\{f\left(W\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{T}\right),V\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{T}\right)\right)|h_{1},\dots,h_{T}\right\}$
$-\mathbb{E}\left\{f\left(W\left(\boldsymbol{Y}_{1},\cdots,\boldsymbol{Y}_{T}\right),V\left(\boldsymbol{Y}_{1},\cdots,\boldsymbol{Y}_{T}\right)\right)|h_{1},\dots,h_{T}\right\}\rightarrow 0.$
@@
Therefore, considering $h_{1},\dots,h_{T}$ as constants. We introduce $\widetilde{W}_{d}=W\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{d-1},\boldsymbol{Y}_{d},\cdots,\boldsymbol{Y}_{T}\right)$ and $\widetilde{V}_{d}=V\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{d-1},\boldsymbol{Y}_{d},\cdots,\boldsymbol{Y}_{T}\right)$ for $d=1,\cdots,T+1,\mathcal{F}_{d}=\sigma\left\{\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{d-1},\boldsymbol{Y}_{d+1},\cdots,\boldsymbol{Y}_{T}\right\}$ for $d=1,\cdots,T$. If there is no danger of confusion, we simply write $\widetilde{W}_{d}$ and $\widetilde{V}_{d}$ as $W_{d}$ and $V_{d}$ respectively (only for this part). Then,
$\left|\mathbb{E}\left\{f\left(W\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{T}\right),V\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{T}\right)\right)\right\}-\mathbb{E}\left\{f\left(W\left(\boldsymbol{Y}_{1},\cdots,\boldsymbol{Y}_{T}\right),V\left(\boldsymbol{Y}_{1},\cdots,\boldsymbol{Y}_{T}\right)\right)\right\}\right|$
$\leq$ $\sum_{d=1}^{T}\mid\mathbb{E}\left\{f\left(W_{d},V_{d}\right)-\mathbb{E}\left\{f\left(W_{d+1},V_{d+1}\right)\right\}\mid\right.$
@@
$V_{d,0}$ $=\beta^{-1}\log\left(\sum_{j=1}^{N}\exp\left(\beta\sqrt{\frac{E(r_{t}^{2})}{T}}\sum_{i=1}^{d-1}\xi_{i,j}+\beta\sqrt{\frac{E(r_{t}^{2})}{T}}\sum_{i=d+1}^{T}Y_{i,j}\right)\right)\in\mathcal{F}_{d}$
@@
Due to $\mathbb{E}\left(\boldsymbol{\xi}_{t}\right)=\mathbb{E}\left(\boldsymbol{Y}_{t}\right)=0$ and $\mathbb{E}\left(\boldsymbol{\xi}_{t}\boldsymbol{\xi}_{t}^{\top}\right)=\mathbb{E}\left(\boldsymbol{Y}_{t}\boldsymbol{Y}_{t}^{\top}\right)$, it can be verified that,
$\mathbb{E}\left(W_{d}-W_{d,0}\mid\mathcal{F}_{d}\right)=\mathbb{E}\left(W_{d+1}-W_{d,0}\mid\mathcal{F}_{d}\right)$
and $\mathbb{E}\left(\left(W_{d}-W_{d,0}\right)^{2}\mid\mathcal{F}_{d}\right)=\mathbb{E}\left(\left(W_{d+1}-W_{d,0}\right)^{2}\mid\mathcal{F}_{d}\right)$. Hence,
$\mathbb{E}\left\{f_{1}\left(W_{d,0},V_{d,0}\right)\left(W_{d}-W_{d,0}\right)\right\}$ $=\mathbb{E}\left\{f_{1}\left(W_{d,0},V_{d,0}\right)\left(W_{d+1}-W_{d,0}\right)\right\}\text{ and }$
$\mathbb{E}\left\{f_{11}\left(W_{d,0},V_{d,0}\right)\left(W_{d}-W_{d,0}\right)^{2}\right\}$ $=\mathbb{E}\left\{f_{11}\left(W_{d,0},V_{d,0}\right)\left(W_{d+1}-W_{d,0}\right)^{2}\right\}.$
Next we consider $V_{d}-V_{d,0}$. Let $z_{d,0,j}=\sqrt{\frac{E(r_{t}^{2})}{T}}\sum_{i=1}^{d-1}\xi_{i,j}+\sqrt{\frac{E(r_{t}^{2})}{T}}\sum_{i=d+1}^{T}Y_{i,j},z_{d,j}=z_{d,0,j}+T^{-1/2}\sqrt{E(r_{t}^{2})}Y_{d,j},z_{d+1,j}=z_{d,0,j}+T^{-1/2}\sqrt{E(r_{t}^{2})}\xi_{d,j}$. By Taylor expansion, we have that:
@@
$=$ $\sum_{l=1}^{T}\partial_{l}F_{\beta}\left(\boldsymbol{z}_{d,0}\right)\left(z_{d,l}-z_{d,0,l}\right)+\frac{1}{2}\sum_{l=1}^{T}\sum_{k=1}^{T}\partial_{k}\partial_{l}F_{\beta}\left(\boldsymbol{z}_{d,0}\right)\left(z_{d,l}-z_{d,0,l}\right)\left(z_{d,k}-z_{d,0,k}\right)$
$+\frac{1}{6}\sum_{l=1}^{T}\sum_{k=1}^{T}\sum_{v=1}^{T}\partial_{v}\partial_{k}\partial_{l}F_{\beta}\left(\boldsymbol{z}_{d,0}+\delta\left(\boldsymbol{z}_{d}-\boldsymbol{z}_{d,0}\right)\right)\left(z_{d,l}-z_{d,0,l}\right)\left(z_{d,k}-\boldsymbol{z}_{d,0,k}\right)\left(\boldsymbol{z}_{d,v}-\boldsymbol{z}_{d,0,v}\right),$
for some $\delta\in(0,1)$. Again, due to $\mathbb{E}\left(\boldsymbol{\xi}_{t}\right)=\mathbb{E}\left(\boldsymbol{Y}_{t}\right)=0$ and $\mathbb{E}\left(\boldsymbol{\xi}_{t}\boldsymbol{\xi}_{t}^{\top}\right)=\mathbb{E}\left(\boldsymbol{Y}_{t}\boldsymbol{Y}_{t}^{\top}\right)$, we can verify that $\mathbb{E}\left\{\left(z_{d,l}-z_{d,0,l}\right)\mid\mathcal{F}_{d}\right\}=\mathbb{E}\left\{\left(z_{d+1,l}-z_{d,0,l}\right)\mid\mathcal{F}_{d}\right\}$ and $\mathbb{E}\left\{\left(z_{d,l}-z_{d,0,l}\right)^{2}\mid\mathcal{F}_{d}\right\}=\mathbb{E}\left\{\left(z_{d+1,l}-z_{d,0,l}\right)^{2}\mid\mathcal{F}_{d}\right)$. By Lemma A.2 in Chernozhukov et al. (2013), we have,
$\left|\sum_{l=1}^{T}\sum_{k=1}^{T}\sum_{v=1}^{T}\partial_{v}\partial_{k}\partial_{l}F_{\beta}\left(\boldsymbol{z}_{d,0}+\delta\left(\boldsymbol{z}_{d}-\boldsymbol{z}_{d,0}\right)\right)\right|\leq C\beta^{2},$
@@
$P\left(\left|\sqrt{E(r_{t}^{2})}\xi_{i,j}\right|\geq t\right)\leq 2\exp\left(-\left(ct/\{\zeta_{1}\sqrt{E(r_{t}^{2})}\}\right)^{\alpha}\right)\lesssim 2\exp\left(-(ct)^{\alpha}\right),$
${P}\left(\max_{1\leq i\leq T}\left|\sqrt{E(r_{t}^{2})}\xi_{ij}\right|>C\log T\right)\rightarrow$ 0 and since $\sqrt{E(r_{t}^{2})}Y_{tj}\sim N(0,1)$,
${P}\left(\max_{1\leq i\leq T}\left|\sqrt{E(r_{t}^{2})}Y_{ij}\right|>C\log T\right)\rightarrow 0.$

### v2

@@ A.4 Proof of Theorem 3
$P\left(\frac{T_{SS}}{\sqrt{2\mbox{tr}({\bf R}^{2})}}\leq x,T||\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}||^{2}_{\infty}\hat{\zeta}-2\log N+\log\log N\leq y\right)\rightarrow\Phi(x)\cdot\exp\left\{-\frac{1}{\sqrt{\pi}}e^{-y/2}\right\}.$ (8)
@@
$P\left(\frac{T_{SS}}{\sqrt{2\mbox{tr}({\bf R}^{2})}}\leq x,T||\hat{{\bf D}}^{-1/2}\hat{\boldsymbol{\theta}}||^{2}_{\infty}\hat{\zeta}\leq u_{N}(y)\right)\rightarrow\Phi(x)\cdot\exp\left\{-\frac{1}{\sqrt{\pi}}e^{-y/2}\right\}.$
@@
$P\left(\sqrt{\frac{1}{2\operatorname{tr}\left(\mathbf{R}^{2}\right)}}N\omega_{T}^{-1}\underset{t_{1}\neq t_{2}}{\sum\sum}h_{t_{1}}h_{t_{2}}{\bf U}_{t_{1}}^{\top}{\bf U}_{t_{2}}+o_{p}\left(1\right)\leq x,\right.$
$\left.\quad N\eta_{\omega}^{-1}\left\|T^{-1/2}\sum_{t=1}^{T}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{t}V_{st}){\bf U}_{t}\right\|_{\infty}^{2}+O_{p}\left(L_{N,T}\right)\leq u_{N}(y)\right)$
@@
where $L_{N,T}=T^{-1/4}\log^{1/2}(TN)+N^{-(1/6\wedge\delta/2)}\log^{1/2}(TN)+T^{-1/2}(\log N)^{1/2}\log^{1/2}(TN)$. Then, we just need to prove (A.4).
First, we will prove that the equation (A.4) holds if ${\bf U}_{t}$ follows the normal distribution. Similar to the proof of Theorem 3 in Feng et al. (2022a), let ${\bf U}_{t}=\left(U_{t,1},\cdots,U_{t,N}\right)^{T}\sim N(\mathbf{0},\boldsymbol{\Sigma}_{u})$. For any set $\Lambda=\left\{i_{1},\cdots,i_{d}\right\}$ with $1\leqslant i_{1}<\cdots<i_{d}\leqslant N$, write $\boldsymbol{U}_{t,\Lambda}=\left({\bf U}_{t,i_{1}},\cdots,{\bf U}_{t,i_{d}}\right)^{T}$ and $\boldsymbol{U}_{t,\Lambda^{C}}=\left({\bf U}_{t,j_{1}},\cdots,{\bf U}_{t,j_{p-d}}\right)^{T}$ where $j_{1}<\cdots<j_{p-d}$ and $\left\{j_{1},\cdots,j_{p-d}\right\}=\{1,2,\dots,N\}\setminus\Lambda$. Define $\mathbf{H}_{t}=\boldsymbol{U}_{t,\Lambda^{C}}-\boldsymbol{\Sigma}_{u,21}\boldsymbol{\Sigma}_{u,11}^{-1}\boldsymbol{U}_{t,\Lambda}$ and $\mathbf{V}_{t}=\boldsymbol{\Sigma}_{u,21}\boldsymbol{\Sigma}_{u,11}^{-1}\boldsymbol{U}_{t,\Lambda}$. Note that
$N\omega_{T}^{-1}\underset{t_{1}\neq t_{2}}{\sum\sum}h_{t_{1}}h_{t_{2}}{\bf U}_{t_{1}}^{\top}{\bf U}_{t_{2}}$
$=$ $\|N^{1/2}\omega_{T}^{-1/2}\sum^{T}_{t=1}h_{t}{\bf U}_{t}\|^{2}-N$
$=$ $\|N^{1/2}\omega_{T}^{-1/2}\sum^{T}_{t=1}h_{t}{\bf U}_{t1,\Lambda}\|^{2}+\|N^{1/2}\omega_{T}^{-1/2}\sum^{T}_{t=1}h_{t}\mathbf{H}_{t}\|^{2}$
$+\|N^{1/2}\omega_{T}^{-1/2}\sum^{T}_{t=1}h_{t}\mathbf{V}_{t}\|^{2}+2N^{1/2}\omega_{T}^{-1/2}\sum^{T}_{t=1}h_{t}\mathbf{V}_{t}^{\top}N^{1/2}\omega_{T}^{-1/2}\sum^{T}_{s=1}h_{s}\mathbf{H}_{s}-N$
and $\omega_{T}^{-1/2}\sum^{T}_{t=1}h_{t}{\bf U}_{t}\sim N(\mathbf{0},\boldsymbol{\Sigma}_{u})$. Hence, according to the proof of Lemma 5.8 in Feng et al. (2022a), under Condition (C5) and $\mbox{tr}(N^{2}\mathbf{\Sigma}_{u}^{2})=\mbox{tr}(\mathbf{R}^{2})\{1+o(1)\}$, for any $\epsilon>0$ there exists $t=t_{N}\rightarrow\infty$ such that
$\max_{\Lambda}P\left(\left|\Theta_{N,\Lambda}\right|\geqslant\epsilon\sqrt{2\mbox{tr}(\mathbf{R}^{2})}\right)\leqslant\frac{1}{N^{t}},$
as $N$ is sufficiently large, where $\Theta_{N,\Lambda}=\|N^{1/2}\omega_{T}^{-1/2}\sum^{T}_{t=1}h_{t}{\bf U}_{t1,\Lambda}\|^{2}+\|N^{1/2}\omega_{T}^{-1/2}\sum^{T}_{t=1}h_{t}\mathbf{V}_{t}\|^{2}+2N^{1/2}\omega_{T}^{-1/2}\sum^{T}_{t=1}h_{t}\mathbf{V}_{t}^{\top}N^{1/2}\omega_{T}^{-1/2}\sum^{T}_{s=1}h_{s}\mathbf{H}_{s}$, the maximum $\Lambda=\left\{i_{1},\cdots,i_{d}\right\}$ runs over all oossible indices $i_{1},\cdots,i_{d}$ with $1\leqslant i_{1}<\cdots<i_{d}\leqslant N$. Hence, similar to the proof of Theorem 3 in Feng et al. (2022a), we can obtain the desired conclusion. We then investigate the non-normal case. Let $\boldsymbol{\xi}_{t}=$ ${\bf U}_{t}\in\mathcal{R}^{N},i=1,2,\cdots,T$. For $\boldsymbol{z}=\left(z_{1},\cdots,z_{q}\right)^{\top}\in\mathcal{R}^{q}$, we consider a smooth approximation of the maximum function, namely,
@@
$V\left(\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{T}\right)$ $=\beta^{-1}\log\left(\sum_{j=1}^{N}\exp\left(\beta\sqrt{\frac{N}{T}}\sum_{i=1}^{T}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{i}V_{si})\boldsymbol{x}_{i,j}\right)\right).$
By setting $\beta$ to satisfy $\log N=o(\beta)$, equation (A.4) is equivalent to
@@
Suppose $\left\{\boldsymbol{Y}_{1},\boldsymbol{Y}_{2},\cdots,\boldsymbol{Y}_{T}\right\}$ are sample from $N\left(0,E{\bf U}_{1}^{\top}{\bf U}_{1}\right)$, and independent with ${\bf U}_{1},\cdots,{\bf U}_{T}$ or write as $\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{T}\right)$. The key idea is to show that: $\left(W\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{T}\right),V\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{T}\right)\right)$ has the same limiting distribution as $\left(W\left(\boldsymbol{Y}_{1},\cdots,\boldsymbol{Y}_{T}\right),V\left(\boldsymbol{Y}_{1},\cdots,\boldsymbol{Y}_{T}\right)\right)$.
@@
Let $l_{b}^{2}(\mathbb{R})$ denote the class of bounded functions with bounded and continuous derivatives up to order 3. It is known that a sequence of randon variables $\left\{Z_{n}\right\}_{n=1}^{\infty}$ converges weakly to a random variable $Z$ if and only if for every $f\in l_{b}^{3}(\mathbb{R}),E\left(f\left(Z_{n}\right)\right)\rightarrow E(f(Z))$. It suffices to show that:
$E\left\{f\left(W\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{T}\right),V\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{T}\right)\right)\right\}-E\left\{f\left(W\left(\boldsymbol{Y}_{1},\cdots,\boldsymbol{Y}_{T}\right),V\left(\boldsymbol{Y}_{1},\cdots,\boldsymbol{Y}_{T}\right)\right)\right\}\rightarrow 0,$
@@
for every $f\in l_{b}^{3}\left(\mathbb{R}^{2}\right)$ as $(N,T)\rightarrow\infty$. We just need to prove
$E\left\{f\left(W\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{T}\right),V\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{T}\right)\right)\right\}$
$-E\left\{f\left(W\left(\boldsymbol{Y}_{1},\cdots,\boldsymbol{Y}_{T}\right),V\left(\boldsymbol{Y}_{1},\cdots,\boldsymbol{Y}_{T}\right)\right)\right\}\rightarrow 0.$
@@
Note that $\{f_{1},\dots,f_{T},r_{1},\dots,r_{T}\}$ are independent with $\boldsymbol{\xi}_{1},\dots,\boldsymbol{\xi}_{T},\boldsymbol{Y}_{1},\dots,\boldsymbol{Y}_{T}$. Therefore, let $\mathcal{F}_{fr}=\sigma\left\{f_{1},\dots,f_{T},r_{1},\dots,r_{T}\right\}$. We introduce $\widetilde{W}_{d}=W\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{d-1},\boldsymbol{Y}_{d},\cdots,\boldsymbol{Y}_{T}\right)$ and $\widetilde{V}_{d}=V\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{d-1},\boldsymbol{Y}_{d},\cdots,\boldsymbol{Y}_{T}\right)$ for $d=1,\cdots,T+1$, $\mathcal{F}_{d}=\sigma\left\{\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{d-1},\boldsymbol{Y}_{d+1},\cdots,\boldsymbol{Y}_{T}\right\}$ for $d=1,\cdots,T$. If there is no danger of confusion, we simply write $\widetilde{W}_{d}$ and $\widetilde{V}_{d}$ as $W_{d}$ and $V_{d}$ respectively (only for this part). Then,
$\left|E\left\{f\left(W\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{T}\right),V\left(\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{T}\right)\right)\right\}-E\left\{f\left(W\left(\boldsymbol{Y}_{1},\cdots,\boldsymbol{Y}_{T}\right),V\left(\boldsymbol{Y}_{1},\cdots,\boldsymbol{Y}_{T}\right)\right)\right\}\right|$
$\leq$ $\sum_{d=1}^{T}\mid E\left\{f\left(W_{d},V_{d}\right)\right\}-E\left\{f\left(W_{d+1},V_{d+1}\right)\right\}\mid$
$=$ $\sum_{d=1}^{T}\mid E\left[E\left\{f\left(W_{d},V_{d}\right)|\mathcal{F}_{fr}\right\}-E\left\{f\left(W_{d+1},V_{d+1}\right)|\mathcal{F}_{fr}\right\}\right]\mid$
$\leq$ $\sum_{d=1}^{T}E\mid E\left\{f\left(W_{d},V_{d}\right)|\mathcal{F}_{fr}\right\}-E\left\{f\left(W_{d+1},V_{d+1}|\mathcal{F}_{fr}\right)\right\}\mid$
@@
$V_{d,0}$ $=\beta^{-1}\log\bigg{(}\sum_{j=1}^{N}\exp\bigg{(}\beta\sqrt{\frac{N}{T}}\sum_{i=1}^{d-1}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{i}V_{si})\xi_{i,j}$
$\quad+\beta\sqrt{\frac{N}{T}}\sum_{i=d+1}^{T}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{i}V_{si})Y_{i,j}\bigg{)}\bigg{)}\in\mathcal{F}_{d}$
@@
Due to $E\left(\boldsymbol{\xi}_{t}\right)=E\left(\boldsymbol{Y}_{t}\right)=0$ and $E\left(\boldsymbol{\xi}_{t}\boldsymbol{\xi}_{t}^{\top}\right)=E\left(\boldsymbol{Y}_{t}\boldsymbol{Y}_{t}^{\top}\right)$, it can be verified that,
$E\left(W_{d}-W_{d,0}\mid\mathcal{F}_{d},\mathcal{F}_{fr}\right)=E\left(W_{d+1}-W_{d,0}\mid\mathcal{F}_{d},\mathcal{F}_{fr}\right)$
and $E\left(\left(W_{d}-W_{d,0}\right)^{2}\mid\mathcal{F}_{d},\mathcal{F}_{fr}\right)=E\left(\left(W_{d+1}-W_{d,0}\right)^{2}\mid\mathcal{F}_{d},\mathcal{F}_{fr}\right)$. Hence,
$E\left\{f_{1}\left(W_{d,0},V_{d,0}\right)\left(W_{d}-W_{d,0}\right)\mid\mathcal{F}_{fr}\right\}$ $=E\left\{f_{1}\left(W_{d,0},V_{d,0}\right)\left(W_{d+1}-W_{d,0}\right)\mid\mathcal{F}_{fr}\right\}\text{ and }$
$E\left\{f_{11}\left(W_{d,0},V_{d,0}\right)\left(W_{d}-W_{d,0}\right)^{2}\mid\mathcal{F}_{fr}\right\}$ $=E\left\{f_{11}\left(W_{d,0},V_{d,0}\right)\left(W_{d+1}-W_{d,0}\right)^{2}\mid\mathcal{F}_{fr}\right\}.$
Next we consider $V_{d}-V_{d,0}$. Let $z_{d,0,j}=\sqrt{\frac{N}{T}}\sum_{i=1}^{d-1}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{i}V_{si})\xi_{i,j}+\sqrt{\frac{N}{T}}\sum_{i=d+1}^{T}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{i}V_{si})Y_{i,j}$, $z_{d,j}=z_{d,0,j}+T^{-1/2}\sqrt{N}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{d}V_{sd})Y_{d,j}$, $z_{d+1,j}=z_{d,0,j}+T^{-1/2}\sqrt{N}(1-\sum_{s=1}^{T}r_{s}^{-1}r_{d}V_{sd})\xi_{d,j}$. By Taylor expansion, we have that:
@@
$=$ $\sum_{l=1}^{N}\partial_{l}F_{\beta}\left(\boldsymbol{z}_{d,0}\right)\left(z_{d,l}-z_{d,0,l}\right)+\frac{1}{2}\sum_{l=1}^{N}\sum_{k=1}^{N}\partial_{k}\partial_{l}F_{\beta}\left(\boldsymbol{z}_{d,0}\right)\left(z_{d,l}-z_{d,0,l}\right)\left(z_{d,k}-z_{d,0,k}\right)$
$+\frac{1}{6}\sum_{l=1}^{N}\sum_{k=1}^{N}\sum_{v=1}^{N}\partial_{v}\partial_{k}\partial_{l}F_{\beta}\left(\boldsymbol{z}_{d,0}+\delta\left(\boldsymbol{z}_{d}-\boldsymbol{z}_{d,0}\right)\right)\left(z_{d,l}-z_{d,0,l}\right)\left(z_{d,k}-\boldsymbol{z}_{d,0,k}\right)\left(\boldsymbol{z}_{d,v}-\boldsymbol{z}_{d,0,v}\right),$
for some $\delta\in(0,1)$. Again, due to $E\left(\boldsymbol{\xi}_{t}\right)=E\left(\boldsymbol{Y}_{t}\right)=0$ and $E\left(\boldsymbol{\xi}_{t}\boldsymbol{\xi}_{t}^{\top}\right)=E\left(\boldsymbol{Y}_{t}\boldsymbol{Y}_{t}^{\top}\right)$, we can verify that $E\left\{\left(z_{d,l}-z_{d,0,l}\right)\mid\mathcal{F}_{d},\mathcal{F}_{fr}\right\}=E\left\{\left(z_{d+1,l}-z_{d,0,l}\right)\mid\mathcal{F}_{d},\mathcal{F}_{fr}\right\}$ and
$E\left\{\left(z_{d,l}-z_{d,0,l}\right)^{2}\mid\mathcal{F}_{d},\mathcal{F}_{fr}\right\}=E\left\{\left(z_{d+1,l}-z_{d,0,l}\right)^{2}\mid\mathcal{F}_{d},\mathcal{F}_{fr}\right\}.$
By Lemma A.2 in Chernozhukov et al. (2013), we have,
$\left|\sum_{l=1}^{N}\sum_{k=1}^{N}\sum_{v=1}^{N}\partial_{v}\partial_{k}\partial_{l}F_{\beta}\left(\boldsymbol{z}_{d,0}+\delta\left(\boldsymbol{z}_{d}-\boldsymbol{z}_{d,0}\right)\right)\right|\leq C\beta^{2},$
@@
$P\left(\left|\sqrt{N}\xi_{i,j}\right|\geq t\right)\leq 2\exp\left(-\left(ct/\{\zeta_{1}\sqrt{N}\}\right)^{\alpha}\right)\lesssim 2\exp\left(-(ct)^{\alpha}\right),$
${P}\left(\max_{1\leq j\leq N}\left|\sqrt{N}\xi_{ij}\right|>C\log N\right)\rightarrow$ 0 and since $\sqrt{N}Y_{tj}\sim N(0,1)$,
${P}\left(\max_{1\leq j\leq N}\left|\sqrt{N}Y_{ij}\right|>C\log N\right)\rightarrow 0.$